### v1

@@ 1 Introduction
This research not only aims to understand the prevalent legal topics within these documents but also seeks to provide insights into the dynamics of legal proceedings and the evolving nature of legal discourse. By uncovering patterns and trends, we can enhance our understanding of legal systems and potentially inform future legal practices and policies. Through this multidimensional analysis, we aim to contribute to the ongoing dialogue surrounding legal document analysis and its implications for the legal profession and society at large.
@@ 2 Related Work
Previous research underscores the significance of legal documents and their widespread implementation. Many of these studies focus on supervised learning techniques utilizing labels. Shukla et al.[7] not only introduce the dataset used in our study but also provide summaries of judgments or segment-wise details, including facts, statutes, and analysis, through various supervised and unsupervised techniques. O et al.[8] concentrate on using topic models to summarize and visualize British legislation to facilitate easier browsing and identification of key legal topics and their associated terms. Wang et al.[9] demonstrate the effectiveness and necessity of experiments to validate decision-making processes in design, highlighting the high performance of LDA algorithm in measuring text similarity. Similarly, Carter et al.[10] conduct similar experiments on legal documents from the High Courts of Australia, focusing on case studies such as the Mabo litigation and the concept of ’proximity’ in tort law. Mohammadi et al.[11] investigate the efficient handling of large-scale legal case law databases like HUDOC, particularly focusing on Article 8 of the European Convention on Human Rights, through topic modeling and citation networks.Kumar et al.[12] propose an approach to generate concise summaries from legal judgments using topics obtained from LDA, providing a notable method for summarization, especially as the first such approach for Indian legal judgments.
Priyadarshini et al.[13] address instability in topic modeling through an ensemble approach, combining Semantic LDA and ensemble models, resulting in reduced processing time compared to conventional methods for legal texts from the UK. Regarding the variety of algorithms for topic modeling, Gonçalves et al.[14] conduct a systematic mapping study to classify and analyze current literature, identifying trends and gaps in research areas and applied methods. Additionally, efforts have been made to enhance the learning of topic models by proposing regularization methods to improve coherence and interpretability, as suggested by NIPS[15].
@@ 3.1 Dataset
The Indian Abstractive dataset (IN-Abs) consists of Indian Supreme Court judgments obtained from the Legal Information Institute of India website, totaling 7,130 case documents with corresponding abstractive summaries. These documents have an average tokens of 5389 in length.
@@
The UK Abstractive dataset (UK-Abs) comprises 793 case documents and their official press summaries from the UK Supreme Court website, segmented into abstractive summaries. These documents have average tokens of 14296 in length.
@@ 4.1 Hyperparameter Tuning
For LDA, hyperparameters are- the number of topics ($k$), $\alpha$ (parameter controlling the sparsity of document-topic distributions), and $\beta$ (parameter controlling the sparsity of topic-word distributions) were meticulously fine-tuned. We tested the model’s performance using various combinations of hyperparameters, including different values of $\alpha$ and $\beta$ ranging from 0.01 to 0.99, both symmetric and asymmetric priors, and a range of $k$ values from 4 to 11. In the India dataset, optimal hyperparameters were determined as $\alpha=0.46$, $\beta=0.91$, and $k=7$, while for the UK dataset, $\alpha$ was set to "asymmetric", $\beta=0.01$, and $k=6$.[20].

### v2

@@ 1 Introduction
This research not only aims to understand the prevalent legal topics within these documents but also seeks to provide insights into the dynamics of legal proceedings and the evolving nature of legal discourse. Uncovering patterns and trends can enhance our understanding of legal systems and inform future legal practices and policies. Through this multidimensional analysis, we aim to contribute to the ongoing dialogue surrounding legal document analysis and its implications for the legal profession and society at large.
@@ 2 Related Work
Previous research underscores the significance of legal documents and their widespread implementation. Many of these studies focus on supervised learning techniques utilizing labels. Shukla et al.[7] not only introduce the dataset used in our study but also provide summaries of judgments or segment-wise details, including facts, statutes, and analysis, through various supervised and unsupervised techniques. O et al.[8] concentrate on using topic models to summarize and visualize British legislation to facilitate easier browsing and identification of key legal topics and their associated terms. Wang et al.[9] demonstrate the effectiveness and necessity of experiments to validate decision-making processes in the design, highlighting the high performance of the LDA algorithm in measuring text similarity. Similarly, Carter et al.[10] conduct similar experiments on legal documents from the High Courts of Australia, focusing on case studies such as the Mabo litigation and the concept of ’proximity’ in tort law. Mohammadi et al.[11] investigate the efficient handling of large-scale legal case law databases like Human Rights Documentation[12], particularly focusing on Article 8 of the European Convention on Human Rights, through topic modeling and citation networks. Kumar et al.[13] propose an approach to generate concise summaries from legal judgments using topics obtained from LDA, providing a notable method for summarization, especially as the first such approach for Indian legal judgments.
Priyadarshini et al.[14] address instability in topic modeling through an ensemble approach, combining Semantic LDA and ensemble models, resulting in reduced processing time compared to conventional methods for legal texts from the UK. Regarding the variety of algorithms for topic modeling, Gonçalves et al.[15] conduct a systematic mapping study to classify and analyze current literature, identifying trends and gaps in research areas and applied methods. Additionally, efforts have been made to enhance the learning of topic models by proposing regularization methods to improve coherence and interpretability, as suggested by Newman et al[16].
@@ 3.1 Dataset
The Indian Abstractive dataset (IN-Abs) consists of Indian Supreme Court judgments obtained from the Legal Information Institute of India website, totaling 7,130 case documents with corresponding abstractive summaries. These documents have an average token, 5389 in length.
@@
The UK Abstractive dataset (UK-Abs) comprises 793 case documents and their official press summaries from the UK Supreme Court website, segmented into abstractive summaries. These documents have average tokens, 14296 in length.
@@ 4.1 Hyperparameter Tuning
For LDA, hyperparameters are- the number of topics ($k$), $\alpha$ (parameter controlling the sparsity of document-topic distributions), and $\beta$ (parameter controlling the sparsity of topic-word distributions) were meticulously fine-tuned. We tested the model’s performance using various combinations of hyperparameters, including different values of $\alpha$ and $\beta$ ranging from 0.01 to 0.99, both symmetric and asymmetric priors, and a range of $k$ values from 4 to 11. In the India dataset, optimal hyperparameters were determined as $\alpha=0.46$, $\beta=0.91$, and $k=7$, while for the UK dataset, $\alpha$ was set to asymmetric, $\beta=0.01$, and $k=6$ [21].